# Machine-precision energy conservative quadrature hyperreduction of Lagrangian hydrodynamics

### v1

@@ Abstract
We present an energy conservative, quadrature based model reduction framework for the compressible Euler equations of Lagrangian hydrodynamics. Building on a high order finite element discretization of the governing equations, we develop a projection based reduced model using data driven reduced basis functions and hyperreduction via the empirical quadrature procedure (EQP). We introduce a strongly energy conservative variant of EQP that enforces exact discrete total energy conservation during the hyperreduction process. Numerical experiments for four benchmark problems—Sedov blast, Gresho vortex, triple point and Taylor–Green vortex—demonstrate that the numerical implementation of our proposed method conserves total energy to near machine precision while maintaining accuracy comparable to the basic EQP formulation. These results establish the energy conservative EQP (CEQP) method as an effective structure preserving hyperreduction strategy for the reduced simulation of nonlinear Lagrangian hydrodynamics.
@@
hydrodynamics, Lagrangian methods, model reduction, hyperreduction, numerical quadrature, energy conservative
@@ 1 Introduction
High fidelity simulations of multiscale, multiphysics systems often involve a large number of degrees of freedom evolving over widely separated temporal and spatial scales, making them prohibitively expensive for tasks such as design, control or uncertainty quantification. Model reduction methods aim to construct low dimensional surrogate models that retain the essential dynamics of the original system while reducing the computational cost associated with its numerical simulation.
For some systems, simplifying physical assumptions can yield reduced models that capture the essential physics of the original system. For complex high dimensional systems such simplifications are often hard to identify. In these cases, data driven methods have emerged as a viable alternative, leveraging simulation or experimental data to construct low dimensional representations that accurately capture the dominant dynamics. In either case, reduction strategies employ a wide variety of mathematical tools including both deterministic and stochastic methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
Despite the different approaches taken in different problem settings, reduction methods can be understood within a single operator theoretic framework which uses projection operators to model the reduction process. In this framework the projected dynamics is governed by equations often referred to as Mori-Zwanzig or Langevin equations, which offer helpful physical interpretations of the different classes of terms that arise in the projection process [13, 14, 15]. In addition to offering guidance about the types of terms that should be approximated in a reduction scheme, this operator theoretic framework can be used to derive error bounds and consistency proofs for the derived reduced models.
@@ 1.1 Projection based reduction
In the present work we use projection based reduction methods to derive reduced models governing the evolution of the projection of the original state variables onto a subspace spanned by a set of reduced basis functions [16].
@@
There are multiple data driven methods for constructing the reduced basis functions used to define a subspace of the state space of the original full dynamics. One of the earliest approaches has been to use the proper orthogonal decomposition (POD) method to construct an orthonormal basis matrix from full simulation data [17, 18, 19]. Given a collection of samples of the state variables, POD yields the best linear fixed-rank approximation to the data in $\mathbb{R}^{N}$ with the euclidean inner product, where $N\in\mathbb{N}$ denotes the state space dimension. An alternative approach is to use the dynamic mode decomposition (DMD) method [20], which can be used to approximate the eigenspaces of the system’s Koopman operator restricted to the finite dimensional subspace spanned by the sampled data [21, 22, 23].
@@
Kernel based methods constitute another class of methods that can be used to derive the reduced basis functions. In kernel methods one employs a symmetric kernel function $k$ to compute the pairwise distances matrix $K_{ij}=k(x_{i},x_{j})$ from state samples $x_{i}$, $x_{j}\in\mathbb{R}^{N}$ and uses the eigenvectors of kernel matrix $K$ as the reduced basis functions. Compared to POD, kernel methods essentially replace the correlation kernel based on the euclidean inner product with the kernel function $k$, thereby embedding the state samples into a reproducing kernel Hilbert space before computing their pairwise distances [24, 25]. Kernel methods can also be used to approximate the eigenspaces of the restricted Koopman operator by using kernels that act on delay embedded state samples [26, 27].
@@
While linear subspace projection has proven successful for reduced models in many applications, its effectiveness is limited to problems where the solution manifold can be accurately approximated by a low dimensional subspace. This assumption is not satisfied in problems with a slow decay in the Kolmogorov $n$-width, such as advection dominated problems that feature sharp gradients, moving shock fronts and turbulence [28]. The presence of such features restricts the effectiveness of reduction based on linear subspace projection. To address this restriction, one proposed alternative has been to create local projection based reduced models by decomposing the solution manifold into submanifolds. Local reduced models are then constructed to approximate the solution within a specific subset of the parameter-time domain during the offline phase. During the online phase, these models are selected based on the parameter, time and current system state. The concept of local reduced models was initially proposed in [29, 30], where unsupervised clustering was employed for the solution manifold decomposition. Windowed reduction approaches were introduced in [31, 32] to build temporally local reduced models that are small but accurate over a short time period for advection dominated problems. These approaches were extended to decomposing the solution manifold based on physical time or other physics based indicators for several problems, including Lagrangian hydrodynamics [33, 34], kinetic plasma [35] and pore collapse in energetic material modeling [36].
@@
All methods outlined above explicitly construct a reduced basis and represent the reduced state variables as a linear combination of the derived basis functions. Another promising approach to alleviate the issue of slow decay in the Kolmogorov $n$-width is to construct a nonlinear projection of the state variables to a subspace of reduced dimension, often called a latent space. The projection map is often represented by a neural network, with the latent space defined by the architecture of the employed neural network and its training procedure [37, 38, 39, 40, 41, 42, 43, 44, 45].
@@
We consider a dynamical system on state space $\mathbb{R}^{N}$, $N\in\mathbb{N}$ governed by the differential equation $\dot{x}=f(x)$ with state variable $x\in X$ and nonlinear function $f:X\to X$. We assume access to an orthonormal reduced basis matrix $\Phi\in\mathbb{R}^{N\times n}$ with $n\ll N$ where each column represents a reduced basis function. Using the representation $x\approx\Phi\hat{x}$, $\hat{x}\in\mathbb{R}^{n}$ we derive the governing equation for the reduced state variable $\hat{x}$ by Galerkin projection
@@
Although this is an equation for the $n\ll N$ reduced state variables, the evaluation of nonlinear function $f$ at every timestep still depends on $\Phi\hat{x}\in\mathbb{R}^{N}$ and is expected to dominate the model’s computational cost. For this reason various hyperreduction methods have been developed to approximate the evaluation of $f$ in different ways.
Hyperreduction methods can be divided into interpolation and quadrature based methods. Interpolation methods employ a set of collocation points to approximate the evaluation of nonlinear function $f$, an idea that can be traced back to [46] in the context of image reconstruction. Examples of this class of methods include the empirical interpolation method [47] and its discrete variant [48]. Interpolation methods employ various sampling strategies to select the collocation points based on different optimality criteria [49, 50, 51, 52]. Quadrature methods offer an alternative strategy for efficiently approximating integrals of nonlinear functions arising in finite element models. They derive reduced quadrature rules that approximate integrations of functions over the computational domain by requiring their evaluation at a limited number of quadrature points. The origin of this approach can be traced back to [53] in the context of computer graphics applications. Examples in this class of methods include the energy conserving sampling and weighting (ECSW) method [54, 55], the empirical cubature method (ECM) [56], the empirical quadrature procedure (EQP) [57, 58, 59] and other methods [60, 61].
In addition to accuracy and reduced complexity, it is often desirable for reduced models to preserve aspects of the qualitative structure of the original full models, such as conserved quantities or symplectic structure. Several reduction schemes have been proposed that incorporate structure preservation constraints directly into the formulation of the reduced model, ensuring that the relevant structure is not destroyed during the employed reduction process [54, 62, 16, 55, 63, 64, 65, 66, 67, 68, 69, 70, 71].
@@ 1.2 Lagrangian hydrodynamics
We consider the compressible Euler equations of fluid dynamics over a three dimensional spatial domain, which can be used to model inviscid, high speed flows and shock wave propagation [72, 73, 74]. The numerical solution of the equations can generally be achieved by two classes of methods: Eulerian and Lagrangian methods. In Eulerian methods the unknown flow variables are treated as fields over a fixed computational mesh, whereas in Lagrangian methods the mesh moves according to the local fluid velocity [75].
Lagrangian methods have traditionally been formulated using approaches such as the staggered grid hydrodynamics (SGH) and the cell centered hydrodynamics (CCH) methods, which usually employ finite difference or finite volume spatial discretizations. The accuracy of these methods relies heavily on the quality of the underlying mesh, which moves and deforms over time and can lead to phenomena such as mesh tangling. One way to address this issue has been the use of arbitrary Lagrangian-Eulerian (ALE) methods which involve a Lagrangian phase and a remesh and remap phase. In the Lagrangian phase, the computational mesh evolves according to the fluid velocity until its quality deteriorates in a prescribed sense. Once that happens, the mesh is adjusted accordingly and the solution fields are remapped onto the adjusted mesh, allowing the Lagrangian simulation to resume. We refer the reader to [72] for details on the methods outlined above and their historical development.
In this work we employ the Lagrangian method developed in [72] which uses a high order finite element discretization of the Euler equations over two- and three-dimensional cartesian grids. Among other features, the employed method offers strong conservation of mass and of discrete total energy. We refer the reader to [72, 76] for the detailed development of the method and its application to various benchmark problems.
1.3 Main contributions
We resume the line of research started in [33, 34], where the authors derived projection based reduced models for Lagrangian hydrodynamics problems using interpolation hyperreduction methods. In the present work we use quadrature hyperreduction methods, which can naturally take advantage of the finite element discretization used in the full model. The main contributions of this work can be summarized as follows.
@@ 1.3 Main contributions
    Projection based reduction. Starting from a finite element (FEM) discretization of the compressible Euler equations, we develop a projection based reduction method using reduced basis functions computed from simulation data. We derive two reduced problem formulations, demonstrating that the reduction process corresponds to replacing the original FEM trial and test functions with the reduced basis functions. We use the empirical quadrature procedure (EQP) to reduce the cost of the computation of integrals of nonlinear force functions that arise in the FEM discretization process.
@@
    Conservative quadrature. We develop an energy conservative variant of the basic EQP method that enforces exact conservation of the discrete total energy in the strong sense in the hyperreduced model. This is achieved by deriving a reduced quadrature rule that preserves the discrete total energy conservation property of the full model.
@@
    Applications. We implement the hyperreduced models derived using the basic and energy conservative EQP methods in the Laghos111 Code repository: https://github.com/CEED/Laghos Lagrangian hydrodynamics simulation code using the libROM222 Code repository: https://github.com/LLNL/libROM library of model reduction methods. We apply the implemented methods to four standard benchmark problems, demonstrating their accuracy and obtained computational speedup. The reported numerical results show that the numerical implementation of the energy conservative EQP method leads to conservation of total energy to near machine precision for all simulated problem cases.
1.4 Paper layout
In Section 2 we present the Euler equations of fluid dynamics and the spatial and temporal discretization methods used to obtain the full model. The projection based reduced model is derived in Section 3, followed by a presentation of the EQP hyperreduction method and its implementation. The energy conservative EQP method is developed in Section 4, where we prove the energy conserving properties of the derived model and outline its numerical implementation. Section 5 presents our numerical results for four different problem cases using both EQP methods, followed by a conclusion in Section 6. A includes additional details on the numerical implementation of the two EQP methods studied in this work.
@@ 2 Full model
We use the isotropic stress tensor $\sigma=-pI+\sigma_{a}$ where $p$ denotes the thermodynamic pressure and $\sigma_{a}$ the artificial viscosity stress. The thermodynamic pressure is governed by an equation of state and can be expressed as a function of the density and internal energy. In the present work we focus on the case of polytropic ideal gases with an adiabatic index $\gamma>1$, which yields the equation of state
@@ 2.1 Spatial discretization
Following [72] we derive a semidiscrete Lagrangian variational formulation which involves a computational mesh that deforms based on the fluid velocity. By the Reynolds transport theorem, the mass conservation in (2.1) is related to the Jacobian $\nabla_{x_{0}}x(t,x_{0})$ of the Lagrangian transformation $x(t,x_{0})$ by
@@
For the spatial discretization of (2.1) we employ a finite element method (FEM) using a kinematic space $\mathcal{V}\subset[H^{1}(\Omega_{0})]^{d}$ with spatial dimension $d\in\mathbb{N}$ and basis $\{\theta_{v,i}\}_{i=0}^{N_{v}-1}$ for approximating the position and velocity fields, and a thermodynamic space $\mathcal{E}\subset L_{2}(\Omega_{0})$ with basis $\{\theta_{e,i}\}_{i=0}^{N_{e}-1}$ for approximating the energy field, where $N_{v}$, $N_{e}\in\mathbb{N}$ denote the number of the global degrees of freedom in the corresponding spaces. In what follows the FEM basis functions are used as both trial and test functions.
@@
$M_{v,ij}=\int_{\Omega(t)}\rho\theta_{i}^{v}\theta_{j}^{v}dx\qquad M_{e,ij}=\int_{\Omega(t)}\rho\theta_{i}^{e}\theta_{j}^{e}$ (2.4)
@@
In the above spatial integrations we write the integrals as taking place over the control volume $\Omega(t)$ instead of $\Omega_{0}=\Omega(0)$. For any scalar valued field $a(t,x_{0})$, $x_{0}\in\Omega_{0}$ this is equivalent to using its Lagrangian extension $a(t,x(t,x_{0}))$, $x(t,x_{0})\in\Omega(t)$ to perform the spatial integration, using the same symbol to denote the two field descriptions. The two approaches are connected by performing the change of variables $x(t,x_{0})\mapsto x_{0}$

### v2

# Machine-precision energy conservative reduced models for Lagrangian hydrodynamics by quadrature methods

@@ Abstract
We present an energy conservative, quadrature based model reduction framework for the compressible Euler equations of Lagrangian hydrodynamics. Building on a finite element discretization of the governing equations, we develop reduced models using data based reduced basis functions and the empirical quadrature procedure (EQP). We introduce a strongly energy conservative variant of EQP that enforces exact energy conservation in the reduction process. Numerical experiments for four benchmark problems—Sedov blast, Gresho vortex, triple point and Taylor–Green vortex—demonstrate that the numerical implementation of our proposed method conserves total energy to near machine precision, while maintaining accuracy comparable to the basic EQP formulation.
@@
model reduction, hyper-reduction, numerical quadrature, energy conservative, hydrodynamics, Lagrangian methods
@@ 1 Introduction
High fidelity simulations of multiscale, multiphysics systems often involve a large number of degrees of freedom evolving over widely separated temporal and spatial scales, making them prohibitively expensive for tasks such as design, control or uncertainty quantification [1, 2]. Model reduction methods—and the related closure or parametrization methods—aim to construct low dimensional surrogate models that retain the essential dynamics of the original system, while reducing the computational cost associated with its numerical simulation [3, 4, 5]. Reduced models can be derived using both physical modeling and data based approaches, depending on the complexity of the underlying dynamics and the availability of high quality data, among other determining factors. In either case, reduction strategies employ a wide variety of mathematical tools, including both deterministic and stochastic methods.
In this work we derive reduced models that govern the evolution of the linear projection of the original state variables onto a subspace spanned by a set of reduced basis functions. Focusing on Lagrangian hydrodynamics, our approach combines physical knowledge about the underlying system in the form of its governing equations, with basis functions computed in a data based fashion.
In the remainder of Section 1 we discuss projection based reduction methods and modeling approaches for Lagrangian hydrodynamics, followed by a summary of the main results of this work. In Section 2 we present the governing equations and the discretization used to obtain the full model. The reduced model is derived in Section 3, followed by a presentation of the EQP hyper-reduction method and its implementation. The energy conservative EQP method is developed in Section 4. Section 5 presents our numerical results for four different problem cases using both EQP methods, followed by a conclusion in Section 6. Finally, A includes additional details on the numerical implementation of the EQP methods.
@@ 1.1 Projection based reduction
Projection based reduction methods derive reduced models by replacing the original state space of the dynamics with a subspace defined as the linear span of a set of reduced basis functions [1, 6]. The reduced state variables are then defined by the linear projection of the original state variables to the constructed subspace. The reduced basis functions that define the state space of the reduced dynamics can be computed by various data based methods, which usually employ the singular value decomposition (SVD) or eigenvalue decomposition of an appropriate data matrix.
@@
The most common method is arguably the proper orthogonal decomposition (POD), which uses the SVD of a snapshot data matrix to construct an orthonormal basis for its range space [7, 8, 9]. Given a collection of samples of the state variables, POD yields the best linear fixed-rank approximation to the data in $\mathbb{R}^{N}$ with the euclidean inner product, where $N\in\mathbb{N}$ denotes the state space dimension. An alternative is the dynamic mode decomposition (DMD) [10], which can be used to approximate the eigenspaces of the system’s Koopman operator restricted to the finite dimensional subspace spanned by the sampled data [11, 12, 13].
@@
The POD and DMD methods can be viewed as specific instances of kernel methods, where one employs a symmetric kernel function $k$ to compute the pairwise distances matrix $K_{ij}=k(u_{i},u_{j})$ from state samples $u_{i}$, $u_{j}\in\mathbb{R}^{N}$, and uses the eigenvectors of the kernel matrix $K$ as the reduced basis functions. Compared to POD, general kernel methods replace the correlation kernel based on the euclidean inner product with the kernel function $k$, thereby embedding the state samples into a feature space before computing their pairwise distances [14]. Kernel methods can also be used to approximate the eigenspaces of the restricted Koopman operator by using kernels that act on delay embedded state samples [15].
@@
The effectiveness of linear subspace projection methods is generally limited to problems where the solution manifold can be accurately approximated by a low dimensional subspace. This assumption is not satisfied in problems with a slow decay in the Kolmogorov $n$-width, such as advection dominated problems that feature sharp gradients, moving shock fronts and turbulence [16, 17]. One approach toward addressing this issue has been to construct local reduced models that approximate the original dynamics for a specific subset of the considered parameter-time domain [18, 19]. A reduction approach based on time windows was introduced in [20, 21] to build temporally local reduced models that are small but accurate over a short time period for advection dominated problems. This approach has since been further extended and applied to problems in various domains [22, 23, 24].
@@
The methods outlined above form a subspace of the original state space and define the reduced state variables by linear projection. An alternative approach is to construct a nonlinear projection of the original state variables onto an appropriate subspace, often called a latent space. The projection map is often represented by a neural network, with the latent space defined by the architecture of the employed neural network and its training procedure [25, 26, 27, 28].
In addition to accuracy and reduced complexity, it is often desirable for reduced models to preserve aspects of the qualitative structure of the original dynamics, such as conserved quantities or symplectic structure. Several reduction schemes have been proposed that incorporate structure preservation constraints directly into the formulation of the reduced model, ensuring that the relevant structure is not destroyed in the reduction process [29, 30, 31, 32, 33, 34, 35, 6, 36, 37, 38, 39, 40].
1.2 Hyper-reduction methods
@@
We consider a dynamical system with state space $\mathbb{R}^{N}$, $N\in\mathbb{N}$, governed by the differential equation $\dot{u}=f(u)$ with state vector $u\in\mathbb{R}^{N}$ and nonlinear function $f\colon\mathbb{R}^{N}\to\mathbb{R}^{N}$. We assume access to an orthonormal reduced basis matrix $\Phi\in\mathbb{R}^{N\times n}$ with $n\ll N$, where each column represents a reduced basis function. Using the representation $u\approx\Phi\hat{u}$, $\hat{u}\in\mathbb{R}^{n}$, we derive the governing equation for the reduced state vector $\hat{u}$ by projection
@@
Although this is an equation for the $n\ll N$ reduced state variables, the evaluation of the nonlinear function $f$ at every timestep depends on $\Phi\hat{u}\in\mathbb{R}^{N}$ and is expected to dominate the model’s computational cost. For this reason, various hyper-reduction methods have been developed to approximate the evaluation of $f$ in different ways.
Hyper-reduction methods can be divided into interpolation and quadrature based methods. Interpolation methods employ a set of collocation points to approximate the evaluation of nonlinear function $f$, an idea that can be traced back to [41] in the context of image reconstruction. Examples of this class of methods include the empirical interpolation method [42] and its discrete variant [43], and employ various sampling strategies to select the collocation points based on different optimality criteria [44, 45, 46, 47]. Quadrature methods offer an alternative strategy for efficiently approximating integrals of nonlinear functions arising in finite element models. They derive reduced quadrature rules that approximate the spatial integration of functions over the computational domain by requiring their evaluation at a limited number of quadrature points. The origin of this approach can be traced back to [48] in the context of computer graphics applications. Examples in this class of methods include the energy conserving sampling and weighting method [29, 31], the empirical cubature method [49], the EQP method [50, 51, 52] and more [53, 54].
@@ 1.3 Lagrangian hydrodynamics
We consider the compressible Euler equations of fluid dynamics in a three dimensional spatial domain, which can be used to model inviscid, high speed flows and shock wave propagation [55, 56, 57]. The numerical solution of the equations can generally be achieved by two classes of methods: Eulerian and Lagrangian methods. In Eulerian methods the unknown variables are treated as fields over a fixed computational mesh, whereas in Lagrangian methods the mesh moves according to the local fluid velocity [58].
Lagrangian methods have traditionally been formulated using approaches such as the staggered grid hydrodynamics and the cell centered hydrodynamics methods, which usually employ finite difference or finite volume spatial discretizations. The accuracy of these methods relies heavily on the quality of the underlying mesh, which moves and deforms over time and can lead to phenomena such as mesh tangling. One way to address this issue has been the use of the arbitrary Lagrangian-Eulerian method, which involves a Lagrangian phase followed by a remesh and remap phase. In the Lagrangian phase, the computational mesh evolves according to the fluid velocity until its quality deteriorates in a prescribed sense. Once that happens, the mesh is adjusted accordingly and the solution fields are remapped onto the adjusted mesh, allowing the Lagrangian simulation to resume. We refer the reader to [57] for details on the methods outlined above and their historical development.
In this work we employ the Lagrangian method developed in [57], which uses a high order finite element discretization of the Euler equations in two and three dimensional cartesian grids. Among other features, the employed method offers strong conservation of mass and discrete total energy. We refer the reader to [57, 59] for the detailed development of the method and its application to various benchmark problems.
1.4 Main results
We resume the line of research initiated in [22, 23], where the authors derived projection based reduced models for Lagrangian hydrodynamics problems using interpolation hyper-reduction methods. In the present work we use quadrature hyper-reduction methods, which are naturally well suited to the finite element discretization used in the full model. The main features of this work can be summarized as follows.
@@ 1.4 Main results
    Projection based reduction. Starting from a finite element (FEM) discretization of the compressible Euler equations, we develop projection based reduction models using reduced basis functions computed from simulation data. We derive two reduced problem formulations, demonstrating that the reduction process corresponds to replacing the original FEM trial and test functions with the reduced basis functions. We use the EQP method to reduce the cost of the spatial integration of nonlinear force functions that arises in the FEM discretization process.
@@
    Conservative quadrature. We develop an energy conservative variant of the basic EQP method that enforces exact conservation of the discrete total energy in the strong sense in the hyper-reduced model. This is achieved by deriving a reduced quadrature rule that preserves the energy conservation property of the full model.
@@
    Applications. We implement the hyper-reduced models derived using the basic and energy conservative EQP methods in the Laghos [^1] Code repository: https://github.com/CEED/Laghos Lagrangian hydrodynamics simulation code using the libROM [^2] Code repository: https://github.com/LLNL/libROM library for model reduction. We apply the methods to four standard benchmark problems, reporting their accuracy and obtained computational speedup. Our numerical results show that the numerical implementation of the energy conservative EQP method leads to conservation of total energy to near machine precision for all simulated problem cases.
@@ 2 Full model
We use the isotropic stress tensor $\sigma=-pI+\sigma_{a}$, where $p$ denotes the thermodynamic pressure and $\sigma_{a}$ the artificial viscosity stress. The thermodynamic pressure is governed by a constitutive equation and can be expressed as a function of the density and internal energy. In the present work we focus on the case of polytropic ideal gases with an adiabatic index $\gamma>1$, which yields the constitutive equation
@@ 2.1 Spatial discretization
We follow [57] to derive a semidiscrete Lagrangian variational formulation based on a computational mesh that deforms according to the fluid velocity. By the Reynolds transport theorem, the mass conservation in (2.1) is related to the Jacobian $\nabla_{x_{0}}x(t,x_{0})$ of the Lagrangian transformation $x(t,x_{0})$ by
@@
For the spatial discretization of (2.1) we employ a finite element method (FEM) using a kinematic space $\mathcal{V}\subset[H^{1}(\Omega_{0})]^{d}$ with spatial dimension $d\in\mathbb{N}$ and basis $\{\theta_{v,i}\}_{i=0}^{N_{v}-1}$ for approximating the position and velocity fields, and a thermodynamic space $\mathcal{E}\subset L_{2}(\Omega_{0})$ with basis $\{\theta_{e,i}\}_{i=0}^{N_{e}-1}$ for approximating the energy field, where $N_{v}$, $N_{e}\in\mathbb{N}$ denote the number of the global degrees of freedom in the corresponding spaces. In what follows, the FEM basis functions are used as both trial and test functions.
@@
$M_{v,ij}=\int_{\Omega(t)}\rho\theta_{i}^{v}\theta_{j}^{v}dx\qquad M_{e,ij}=\int_{\Omega(t)}\rho\theta_{i}^{e}\theta_{j}^{e}dx$ (2.4)
@@
In the above spatial integrations we write the integrals as taking place over the control volume $\Omega(t)$ instead of $\Omega_{0}=\Omega(0)$. For any scalar valued field $a(t,x_{0})$, $x_{0}\in\Omega_{0}$, this is equivalent to using its Lagrangian extension $a(t,x(t,x_{0}))$, $x(t,x_{0})\in\Omega(t)$, to perform the spatial integration; without misunderstanding, we use the same symbol to denote the two field descriptions. The two approaches are connected by performing the change of variables $x(t,x_{0})\mapsto x_{0}$